A student titrates a 20.13 mL sample of HCl of unknown concentration. She begins her titration at the 0.00 mL mark and ends the titration at the 33.98 mL mark. The NaOH solution has a molarity of 1.232. What is the concentration of HCl?
1
Expert's answer
2020-10-16T08:56:53-0400

HCl + NaOH = NaCl + H2O

NHCl * Vsolution HCl = NNaOH * Vsolution NaOH

N = C/feq, feq = 1, N=C.

CHCl * Vsolution HCl = CNaOH * Vsolution NaOH

CHCl = (CNaOH * Vsolution NaOH)/Vsolution HCl

Vsolution NaOH = 33,98 mL = 33,98 * 10-3 L

Vsolution HCl = 20,13 ml = 20,13 * 10-3 L

CHCl = (1,232 * 33,98 *10-3)/(20,13 * 10-3) = 2,080.